# On the power of graph neural networks and the role of the activation function

### v4

@@ Abstract.
In this article we present new results about the expressivity of Graph Neural Networks (GNNs). We prove that for any GNN with piecewise polynomial activations, whose architecture size does not grow with the graph input sizes, there exists a pair of non-isomorphic rooted trees of depth two such that the GNN cannot distinguish their root vertex up to an arbitrary number of iterations. The proof relies on tools from the algebra of symmetric polynomials. In contrast, it was already known that unbounded GNNs (those whose size is allowed to change with the graph sizes) with piecewise polynomial activations can distinguish these vertices in only two iterations. It was also known prior to our work that with ReLU (piecewise linear) activations, bounded GNNs are weaker than unbounded GNNs [${}^{+}$22]. Our approach adds to this result by extending it to handle any piecewise polynomial activation function, which goes towards answering an open question formulated by Grohe [Gro21] more completely. Our second result states that if one allows activations that are not piecewise polynomial, then in two iterations a single neuron perceptron can distinguish the root vertices of any pair of nonisomorphic trees of depth two (our results hold for activations like the sigmoid, hyperbolic tan and others). This shows how the power of graph neural networks can change drastically if one changes the activation function of the neural networks. The proof of this result utilizes the Lindemann-Weierstrauss theorem from transcendental number theory.
@@ 1. Introduction
Graph Neural Networks (GNNs) form a popular framework for a variety of computational tasks involving network data, with applications ranging from analysis of social networks, structure and functionality of molecules in chemistry and biological applications, computational linguistics, simulations of physical systems, techniques to enhance optimization algorithms, to name a few. The interested reader can look at [ZAL18, ${}^{+}$17, ${}^{+}$19, DBV16, ${}^{+}$08, Ham20, ${}^{+}$15, ${}^{+}$17, ${}^{+}$20, ${}^{+}$20, ${}^{+}$16, ${}^{+}$21, ${}^{+}$20], which is a small sample of a large and actively growing body of work.
@@
At a high level, the computational models of GNNs, Wesfeiler-Lehman/color refinement type algorithms and certain logics in descriptive complexity are intimately connected because they all fall under the paradigm of trying to discern something about the global structure of a graph from local neighborhood computations. Informally, these algorithms iteratively maintain a state (a.k.a. “color”) for each vertex of the graph and in every iteration, the state of a vertex is updated by performing some predetermined set of operations on the set of current states of its neighbors (including itself). The different kinds of allowed states and allowed operations determine the computational paradigm. For instance, in GNNs, the states are typically vectors in some Euclidean space and the operations for updating the state are functions that can be represented by deep neural networks. As another example, in the color refinement algorithm, the states are multisets of some predetermined finite class of labels and the operations are set operations on these multisets. A natural question then arises: Given two of these models, which one is more powerful, or equivalently, can one of the models always simulate the other? Several mathematically precise answers to such questions have already been obtained. For instance, it has been proved independently by [${}^{+}$19] and [XHLJ18] that the color refinement algorithm precisely captures the expressiveness of GNNs in the sense that there is a GNN distinguishing two nodes of a graph (by assigning them different state vectors) if and only if color refinement assigns different multisets to these nodes. Such a characterization holds for unbounded GNNs, i.e. GNNs for which the underlying neural network sizes can grow with the size of the input graphs. This implies a characterization of the distinguishability of nodes by GNNs as being equivalent to what is known as Graded Modal Counting Logic (GC2); see [${}^{+}$20] for some recent, quantitatively precise results in this direction.
@@
Reviewing these equivalences in a recent survey [Gro21], Grohe emphasizes the fact that the above mentioned equivalence between the separation power of GNNs and the color refinement algorithm has only been established for unbounded GNNs whose neural network sizes are allowed to grow as a function of the size of the input graphs. Question 1 on his list of important open questions in this topic asks what happens if one considers bounded GNNs, i.e., the the size of the neural networks is fixed a priori and cannot change as a function of the size of the input graphs. Do bounded GNNs have the same separation power as unbounded GNNs and color refinement? In [${}^{+}$22], the authors provide a negative answer to this question in the case of Rectified Linear Unit (ReLU) activations, by proving lower bounds on the size of the GNNs with ReLU activations that are needed to simulate color refinement. Their approach combines results from communication complexity and properties of ReLU neural networks, including upper bounds on the number of their linear regions. In this article, we answer this question in the negative in the more general case of piecewise polynomial activations. Given any bounded GNN with such activations, we explicitly construct two non isomorphic rooted trees of depth two such that their root nodes cannot be distinguished by the GNN. Interestingly, only the sizes of the trees depend on the GNN, but their depth does not. This result is stated formally in Theorem 3 and it holds for bounded GNNs with piecewise polynomial activations (this includes, e.g., ReLU activations). We prove a second result that shows how the activation function dramatically impacts the expressivity of bounded size GNNs: if one allows activation functions that are not piecewise polynomial, all root nodes of rooted trees of depth two can be distinguished by a single neuron perceptron. This result is formally stated in Theorem 4. In spite of this, it is not clear whether these more general activations allow one to simulate color refinement with bounded GNNs for all graphs. Thus, there remains future work to be done in answering the above question from Grohe’s survey fully.
@@ 2. Formal statement of results
We assume graphs to be finite, undirected, simple, and vertex-labelled: a graph is a tuple $G=(V(G),E(G),P_{1}(G),...,P_{\ell}(G))$ consisting of a finite vertex set $V(G)$, a binary edge relation $E(G)\subset V(G)^{2}$ that is symmetric and irreflexive, and unary relations $P_{1}(G),\cdots,P_{\ell}(G)\subset V(G)$ representing $\ell>0$ vertex labels. In the following, we suppose that the $P_{i}(G)$’s form a partition of the set of vertices of $G$, i.e. each vertex has a unique label. Also, the number $\ell$ of labels, which we will also call colors, is supposed to be fixed and does not grow with the size of the input graphs. When there is no ambiguity about which graph $G$ is being considered, $N(v)$ refers to the set of neighbors of $v$ in $G$ not including $v$. $\lvert G\rvert$ will denote the number of vertices of $G$. We use simple curly brackets for a set $X=\{x\in X\}$ and double curly brackets for a multiset $Y=\{\{y\in Y\}\}$. For a set $X$, $\lvert X\rvert$ is the cardinal of $X$. When $m$ is a positive integer, $\mathfrak{S}_{m}$ is the set of permutations of $\{1,\cdots,m\}$.
@@
Let $m$ be a positive integer. A function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is piecewise polynomial iff there exist multivariate polynomials $P_{1},\cdots,P_{r}\in\mathbb{R}[X_{1},\cdots,X_{m}]$ such that for any $x\in\mathbb{R}^{m}$, there exists $i\in\{1,\cdots,r\}$ such that $f(x)=P_{i}(x)$. The degree of a piecewise polynomial function $f$ is $\mathsf{deg}(f):=\max\{\mathsf{deg}(P_{1}),\cdots,\mathsf{deg}(P_{r})\}$. The number of polynomial pieces of a piecewise polynomial $f$ is the smallest $r$ such that $f$ can be represented as above.
Definition 2 ((Elementary) Symmetric polynomial).
For any positive integer $m$, a polynomial $P\in\mathbb{R}[X_{1},\cdots,X_{m}]$ is said to be symmetric if for any permutation $\pi\in\mathfrak{S}_{m}$ of $\{1,\cdots,m\}$ and any $v_{1},\ldots,v_{m}\in\mathbb{R}$, $P(v_{\pi(1)},\cdots,v_{\pi(m)})=P(v_{1},\cdots,v_{m})$. For any $k\in\{1,\cdots,m\}$, the elementary symmetric polynomial $s_{k}$ is given by $s_{k}(X_{1},\cdots,X_{m}):=\sum_{1\leq j_{1}<j_{2}<\cdots<j_{k}\leq m}X_{j_{1}}\dotsm X_{j_{k}}$.
@@
[Gro21, XHLJ18, ${}^{+}$19] Let $d\geq 1$, and let $\xi^{d}$ be an embedding computed by a GNN after $d$ iterations. Then $\mathsf{cr}^{d}$ refines $\xi$, that is, for all graphs $G,G^{\prime}$ and vertices $v\in V(G)$, $v^{\prime}\in V(G^{\prime})$, $\mathsf{cr}^{d}(v)=\mathsf{cr}^{d}(v^{\prime})\implies\xi^{d}(G,v)=\xi^{d}(G^{\prime},v^{\prime})$.
@@
[Gro21, Theorem VIII.4][XHLJ18, ${}^{+}$19] Let $n\in\mathbb{N}$. Then there is a recurrent GNN such that for all $t=0,\cdots,n$, the vertex invariant $\xi^{t}$ computed in the $t$-th iteration of the GNN refines $\mathsf{cr}^{t}$ on all graphs of order at most $n$.
@@
Remark 3.
Furthermore, our proof also shows that for a GNN to distinguish the source vertices of any pair of trees of depth $2$ (displayed in Figure 1) with $m$ vertices at depth one and with prescribed degrees $k_{1},\cdots,k_{m}\in\{1,\cdots M\}$, it is necessary that the number of regions $r$ of the underlying neural network of degree $q$ satisfies the following inequality:
(1) $r\geq\binom{M+m+1}{m}\frac{1}{M^{q^{2}}(\binom{m}{q}q)^{q}}$
This condition may not be sufficient, as the polynomial regions have also to be correctly placed. Nevertheless, for ReLU neural networks ($q=1$), there exist upper bounds for the number of regions given the number of layers and neurons per each layer (cf. [MPCB14, Mon17] and [STR18] for refinements). Therefore, for a ReLU neural network with $L$ layers, input dimension $n_{0}$ and with $n_{1},\cdots,n_{L}$ neurons per layer, it is necessary that
$\prod_{l=1}^{L-1}\Bigl{\lfloor}\frac{n_{l}}{n_{0}}\Bigr{\rfloor}^{n_{0}}\cdot\sum_{j=0}^{n_{0}}\binom{n_{L}}{j}\geq\binom{M+m+1}{m}\frac{1}{M\cdot m}$
This can be used to obtain quantitative lower bounds on the size of the neural network. One could also derive similar lower bounds in the general piecewise polynomial case if provided an upper bound on the number of polynomial regions given the number of layers and neurons per layer.
@@ 3. Overview of the proofs
The proof of the first result is structured as follows. Since the trees are parameterized by $m$-tuples of integers $k_{1},\ldots,k_{m}$, the embedding of the root node computed by the GNN at any iteration is a function of these $m$ integers. Since the activations are piecewise polynomials, these embeddings of the root node are also piecewise multivariate symmetric polynomial functions of $k_{1},\ldots,k_{m}$ (Lemma 3). Then, we show that there exists a large enough region of $\mathbb{R}^{m}$ on which this piecewise polynomial function is evaluated by the same polynomial. This region is large enough in the following sense: we prove that it contains more integral vectors than the number of possible values a symmetric polynomial with bounded degree can take on these vectors, even after identifying vectors up to permutations of the coordinates. This implies that the polynomial will take the same value on two distinct integral vectors whose coordinates are not identical up to permutations. When translating this result back to the world of GNNs, this implies that the two embeddings of the root nodes of the trees corresponding to these two vectors will coincide. To conclude a separation between bounded and unbounded GNNs, we justify that the unbounded ones can seperate these two vertices. This is based on the previous result (Theorem 2) stating that unbounded GNNs refine color refinement.
@@ 4. Collision with piecewise polynomial activations
The following statement is a reformulation of the fundamental theorem of symmetric polynomials, where we add a constraint on the degree of the decomposition. We provide a proof for completeness.
Lemma 1.
Let $P$ be a symmetric multivariate polynomial of $m$ variables of degree $q$ with $q\leq m$. Then $P$ can be written as a polynomial of degree $q$ of the elementary symmetric polynomials $s_{1},\cdots,s_{q}$.
Proof.
For $\alpha=(\alpha_{1},\cdots,\alpha_{m})\in\mathbb{N}^{m}$, define the multidegree of a monomial:
$\mathsf{mdeg}(X_{1}^{\alpha_{1}}\cdots X_{m}^{\alpha_{m}}):=\sum_{i=1}^{m}\alpha_{i}(q+1)^{m-i}.$
By definition, the leading term of a polynomial is the monomial with greatest multidegree. We present a proof by induction on the multidegree of $P$. We first need the following claim.
Claim: Let $P\in\mathbb{R}[X_{1},\cdots,X_{m}]$ be a symmetric polynomial. Let $c_{\alpha}X^{\alpha}=c_{\alpha}X_{1}^{\alpha_{1}}\cdots X_{m}^{\alpha_{m}}$ be a monomial of $P$. Then $c_{\alpha}X^{\pi(\alpha)}$ is also a monomial in $P$, for every permutation $\pi$ of $\{1,\cdots,m\}$.
Proof of the claim.
Define $\phi^{\pi}:\mathbb{R}[X_{1},\cdots,X_{m}]\rightarrow\mathbb{R}[X_{1},\cdots X_{m}]$, be the linear map between polynomials which maps variable $X_{i}$ to $X_{\pi(i)}$. Since $P$ is symmetric, then $\phi^{\pi}(P)=P$ (we use here the equivalence between equality of $\phi^{\pi}(P)$ and $P$ as functions and as formal algebraic objects). Since $c_{\alpha}X^{\alpha}$ is a monomial in $P$ and $c_{\alpha}X^{\pi(\alpha)}$ is a monomial in $\phi^{\pi}(P)$ for any permutation $\pi$, we must have $c_{\alpha}X^{\pi(\alpha)}$ as a monomial in $P$. ∎
Base case. The property is true for any polynomial of multidegree 00 (constant polynomial).
Induction step. Let $c_{\alpha}X^{\alpha}$ be the leading term of $P$. Then $c_{\alpha}X^{\pi(\alpha)}$ is a monomial in $P$ for every $\pi\in\mathfrak{S}_{m}$ by the previous claim. Therefore, the leading term’s exponent $\alpha=(\alpha_{1},\cdots,\alpha_{m})$ must satisfy $\alpha_{1}\geq\alpha_{2}\geq\cdots\geq\alpha_{m}$. Since $P$ has degree $q$ then $\alpha_{i}=0$ for any $i\geq q+1$.
Let $d_{m}=\alpha_{m}$, $d_{m-1}=\alpha_{m-1}-\alpha_{m},\cdots,d_{i}=\alpha_{i}-\alpha_{i+1},\cdots,d_{1}=\alpha_{1}-\alpha_{2}$. Define $Q^{\prime}(X_{1},\cdots,X_{m}):=s_{1}^{d_{1}}\cdots s_{m}^{d_{m}}$. $Q^{\prime}$ is a polynomial of $s_{1},\cdots,s_{q}$ and the leading term of $Q^{\prime}$ is $c^{\prime}_{\alpha}X^{\alpha}$ where $c^{\prime}_{\alpha}\neq 0$. In particular, $\mathsf{deg}(Q^{\prime})\leq q$.
Now, let $P^{\prime}:=P-\frac{c_{\alpha}}{c^{\prime}_{\alpha}}Q^{\prime}$. Then $\mathsf{mdeg}(P^{\prime})<\mathsf{mdeg}(P)$, and $P^{\prime}$ is symmetric because $P$ and $Q^{\prime}$ are. Applying the induction hypothesis to $P^{\prime}$, get $Q^{\prime\prime}$ such that $P^{\prime}=Q^{\prime\prime}(s_{1},\cdots,s_{q})$. Define $Q:=\frac{c_{\alpha}}{c^{\prime}_{\alpha}}Q^{\prime}+Q^{\prime\prime}$. $Q$ is a polynomial of degree at most $q$ of $s_{1},\cdots,s_{q}$ and $Q=P$, which completes the induction step. ∎
@@
On the other hand, $\lvert\mathsf{Dom}(\Phi)\rvert=\binom{M+m-1}{m}$ (number of multisets of size $m$ whose elements are taken from a set of size $M$). Now, let $f_{m}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ be the $m$-th function in the given sequence of symmetric piecewise polynomial functions where each polynomial piece has degree at most $q$. Let $r$ be the number of pieces of $f_{m}$. Then, there is a subset of $\mathsf{Dom}(\Phi)$ with at least $\frac{1}{r}\binom{M+m-1}{m}$ elements where $f_{m}$ is a symmetric polynomial $P$ of degree at most $q$. Lemma 1 tells us that $P$ can be expressed as a polynomial of degree at most $q$ of the elementary symmetric polynomials $s_{1},\cdots,s_{q}$. Due to the pigeonhole principle, any such polynomial will be equal on two distinct multisets $\{\{k_{1},\cdots,k_{m}\}\}$ and $\{\{k^{\prime}_{1},\cdots,k^{\prime}_{m}\}\}$ in $\Omega_{M}$ as soon as:
@@
Let $\xi^{t}(T[k_{1},\ldots,k_{m}],s)$ be the embedding obtained via a GNN with piecewise activation functions after $t$ iterations, where $\xi^{0}(v)=1$ for all vertices $v\in V(T[k_{1},\ldots,k_{m}])$. Then, for any iteration $t$, there exists a symmetric multivariate piecewise polynomial function $F$ such that $\xi^{t}(T[k_{1},\ldots,k_{m}],s)=F(k_{1},\cdots,k_{m})$. Furthermore, the degree of $F$ does not depend on $m$, but only on the underlying neural network and $t$.
@@
We first prove by induction on $t$ that, for any vertex $v\in V(T[k_{1},\cdots,k_{m}])$, $\xi^{t}(T[k_{1},\ldots,k_{m}],v)$ is a piecewise polynomial function of the $k_{i}$’s.
Base case: for $t=0$ this is trivial since all vertices are initialised with the constant polynomial 1, whose degree does not depend on $m$.
Induction step: Suppose the property is true at iteration $t$, i.e for each node $w$, $\xi^{t}(T[k_{1},\ldots,k_{m}],w)$ is a multivariate polynomial of the $k_{i}$’s. Since

### v5

@@ Abstract.
In this article we present new results about the expressivity of Graph Neural Networks (GNNs). We prove that for any GNN with piecewise polynomial activations, whose architecture size does not grow with the graph input sizes, there exists a pair of non-isomorphic rooted trees of depth two such that the GNN cannot distinguish their root vertex up to an arbitrary number of iterations. In contrast, it was already known that unbounded GNNs (those whose size is allowed to change with the graph sizes) with piecewise polynomial activations can distinguish these vertices in only two iterations. It was also known prior to our work that with ReLU (piecewise linear) activations, bounded GNNs are weaker than unbounded GNNs [ACI+22]. Our approach adds to this result by extending it to handle any piecewise polynomial activation function, which goes towards answering an open question formulated by Grohe [Gro21] more completely. Our second result states that if one allows activations that are not piecewise polynomial, then in two iterations a single neuron perceptron can distinguish the root vertices of any pair of nonisomorphic trees of depth two (our results hold for activations like the sigmoid, hyperbolic tan and others). This shows how the power of graph neural networks can change drastically if one changes the activation function of the neural networks. The proof of this result utilizes the Lindemann-Weierstrauss theorem from transcendental number theory.
@@ 1. Introduction
Graph Neural Networks (GNNs) form a popular framework for a variety of computational tasks involving network data, with applications ranging from analysis of social networks, structure and functionality of molecules in chemistry and biological applications, computational linguistics, simulations of physical systems, techniques to enhance optimization algorithms, to name a few. The interested reader can look at [ZAL18, BBL+17, DZC+19, DBV16, SGT+08, Ham20, DMI+15, KDZ+17, SYS+20, ZCH+20, BPL+16, CCK+21, SGGP+20], which is a small sample of a large and actively growing body of work.
@@
At a high level, the computational models of GNNs, Wesfeiler-Lehman/color refinement type algorithms and certain logics in descriptive complexity are intimately connected because they all fall under the paradigm of trying to discern something about the global structure of a graph from local neighborhood computations. Informally, these algorithms iteratively maintain a state (a.k.a. “color”) for each vertex of the graph and in every iteration, the state of a vertex is updated by performing some predetermined set of operations on the set of current states of its neighbors (including itself). The different kinds of allowed states and allowed operations determine the computational paradigm. For instance, in GNNs, the states are typically vectors in some Euclidean space and the operations for updating the state are functions that can be represented by deep neural networks. As another example, in the color refinement algorithm, the states are multisets of some predetermined finite class of labels and the operations are set operations on these multisets. A natural question then arises: Given two of these models, which one is more powerful, or equivalently, can one of the models always simulate the other? Several mathematically precise answers to such questions have already been obtained. For instance, it has been proved independently by [MRF+19] and [XHLJ18] that the color refinement algorithm precisely captures the expressiveness of GNNs in the sense that there is a GNN distinguishing two nodes of a graph (by assigning them different state vectors) if and only if color refinement assigns different multisets to these nodes. Such a characterization holds for unbounded GNNs, i.e. GNNs for which the underlying neural network sizes can grow with the size of the input graphs. This implies a characterization of the distinguishability of nodes by GNNs as being equivalent to what is known as Graded Modal Counting Logic (GC2); see [BKM+20] for some recent, quantitatively precise results in this direction.
@@
Reviewing these equivalences in a recent survey [Gro21], Grohe emphasizes the fact that the above mentioned equivalence between the separation power of GNNs and the color refinement algorithm has only been established for unbounded GNNs whose neural network sizes are allowed to grow as a function of the size of the input graphs. Question 1 on his list of important open questions in this topic asks what happens if one considers bounded GNNs, i.e., the the size of the neural networks is fixed a priori and cannot change as a function of the size of the input graphs. Do bounded GNNs have the same separation power as unbounded GNNs and color refinement? In [ACI+22], the authors provide a negative answer to this question in the case of Rectified Linear Unit (ReLU) activations, by proving lower bounds on the size of the GNNs with ReLU activations that are needed to simulate color refinement. Their approach combines results from communication complexity and properties of ReLU neural networks, including upper bounds on the number of their linear regions. In this article, we answer this question in the negative in the more general case of piecewise polynomial activations. Given any bounded GNN with such activations, we explicitly construct two non isomorphic rooted trees of depth two such that their root nodes cannot be distinguished by the GNN. Interestingly, only the sizes of the trees depend on the GNN, but their depth does not. This result is stated formally in Theorem 3 and it holds for bounded GNNs with piecewise polynomial activations (this includes, e.g., ReLU activations). We prove a second result that shows how the activation function dramatically impacts the expressivity of bounded size GNNs: if one allows activation functions that are not piecewise polynomial, all root nodes of rooted trees of depth two can be distinguished by a single neuron perceptron. This result is formally stated in Theorem 4. [AGA+24] provide a more general result for analytic non polynomial activation functions: it is possible to simulate color refinement on any graph if one allows analytic non polynomial functions. [BKR24] also showed that among GNNs with analytic non polynomial activations, those with one dimensional embeddings at every iteration are sufficient to simulate color-refinement. In contrast with approaches presented in [AGA+24, BKR24], our proof does not use the analytic properties of the activation function, but linear independence over the integers of the activation function evaluated at distinct integers.
@@ 2. Formal statement of results
We assume graphs to be finite, undirected, simple, and vertex-labelled: a graph is a tuple $G=(V(G),E(G),P_{1}(G),...,P_{\ell}(G))$ consisting of a finite vertex set $V(G)$, a binary edge relation $E(G)\subset V(G)^{2}$ that is symmetric and irreflexive, and unary relations $P_{1}(G),\cdots,P_{\ell}(G)\subset V(G)$ representing $\ell>0$ vertex labels. In the following, we suppose that the $P_{i}(G)$’s form a partition of the set of vertices of $G$, i.e. each vertex has a unique label. Also, the number $\ell$ of labels, which we will also call colors, is supposed to be fixed and does not grow with the size of the input graphs. When there is no ambiguity about which graph $G$ is being considered, $N(v)$ refers to the set of neighbors of $v$ in $G$ not including $v$. $\lvert G\rvert$ will denote the number of vertices of $G$. We use simple curly brackets for a set $X=\{x\in X\}$ and double curly brackets for a multiset $Y=\{\{y\in Y\}\}$. For a set $X$, $\lvert X\rvert$ is the cardinal of $X$. When $m$ is a positive integer, $\mathfrak{S}_{m}$ is the set of permutations of $\{1,\cdots,m\}$. We will also need the following basic notions from algebra. Given any ring $(R,+,\times)$, $R[X_{1},\ldots,X_{m}]$ will denote the polynomial ring in $m$ indeterminates with coefficients in $R$. An (additive) subgroup generated by $\sigma_{1},\cdots,\sigma_{q}\in R$ is the smallest (inclusion-wise) additive subgroup of $R$ that contains $\sigma_{1},\cdots,\sigma_{q}$. Equivalently, it corresponds to the set $\{\lambda_{1}\sigma_{1}+\cdots+\lambda_{q}\sigma_{q}:\lambda\in\mathbb{Z}^{q}\}$.
@@
Let $m$ be a positive integer. A function $f:\mathbb{R}^{m}\rightarrow\mathbb{R}$ is piecewise polynomial iff there exist multivariate polynomials $P_{1},\cdots,P_{r}\in\mathbb{R}[X_{1},\cdots,X_{m}]$ such that for any $x\in\mathbb{R}^{m}$, there exists $i\in\{1,\cdots,r\}$ such that $f(x)=P_{i}(x)$. A polynomial region $A$ of $f$ is a set such that there exists $i\in[r]$ such that $A\subseteq\{x\in\mathbb{R}^{m}:f(x)=P_{i}(x)\}$. The degree of a piecewise polynomial function $f$ is $\mathsf{deg}(f):=\max\{\mathsf{deg}(P_{1}),\cdots,\mathsf{deg}(P_{r})\}$. The number of polynomial pieces of a piecewise polynomial $f$ is the smallest $r$ such that $f$ can be represented as above.
Definition 2 (Finitely generated polynomial).
Let $q$ be a positive integer. A multivariate polynomial $P$ is $q$-generated provided there exist reals $\sigma_{1},\cdots,\sigma_{q}$ such that $P=\sum_{\alpha\in S}\gamma_{\alpha}X^{\alpha}$, where $\gamma_{\alpha}$ belongs to the additive subgroup of $\mathbb{R}$ generated by $\sigma_{1},\cdots,\sigma_{q}$, and $S$ is the set of exponents of $P$. Under these conditions, we say that $P$ is $q$-generated by the reals $\sigma_{1},\cdots,\sigma_{q}$.
@@
[Gro21, XHLJ18, MRF+19] Let $d\geq 1$, and let $\xi^{d}$ be an embedding computed by a GNN after $d$ iterations. Then $\mathsf{cr}^{d}$ refines $\xi$, that is, for all graphs $G,G^{\prime}$ and vertices $v\in V(G)$, $v^{\prime}\in V(G^{\prime})$, $\mathsf{cr}^{d}(v)=\mathsf{cr}^{d}(v^{\prime})\implies\xi^{d}(G,v)=\xi^{d}(G^{\prime},v^{\prime})$.
@@
[Gro21, Theorem VIII.4][XHLJ18, MRF+19] Let $n\in\mathbb{N}$. Then there is a recurrent GNN such that for all $t=0,\cdots,n$, the vertex invariant $\xi^{t}$ computed in the $t$-th iteration of the GNN refines $\mathsf{cr}^{t}$ on all graphs of order at most $n$.
@@ 3. Overview of the proofs
The proof of the first result is structured as follows. Since the trees are parameterized by $m$-tuples of integers $k_{1},\ldots,k_{m}$, the embedding of the root node computed by the GNN at any iteration is a function of these $m$ integers. Since the activations are piecewise polynomials, these embeddings of the root node are also piecewise multivariate polynomial functions of $k_{1},\ldots,k_{m}$. We further prove that the complexity of its polynomial pieces can be controlled uniformly in the sense of Definition 2: the number of generators of such polynomial is independent of $m$, but depends only on the number of iterations of the GNN and the underlying neural network (Lemma 2 and 3). Then, we show that there exists a large enough region of $\mathbb{R}^{m}$ on which this piecewise polynomial function is evaluated by the same polynomial. This region is large enough in the following sense: we prove that for a dimension $m$ that is sufficiently large, the region contains more integral vectors than the number of possible values a $q$-generated polynomial with degree at most $q$ can take on these vectors, even after identifying vectors up to permutations of the coordinates (Lemmas LABEL:lemmanbvaluesgeneratedp and 1). This implies that the polynomial piece will take the same value on two distinct integral vectors whose coordinates are not identical up to permutations. When translating this result back to the world of GNNs, this means that the two embeddings of the root nodes of the trees corresponding to these two vectors will coincide. To conclude a separation between bounded and unbounded GNNs, we justify that the unbounded ones can seperate these two vertices. This is based on the previous result (Theorem 2) stating that unbounded GNNs refine color refinement.
@@ 4. Collision with piecewise polynomial activations
On the other hand, $\lvert\mathsf{Dom}(\Phi)\rvert=\binom{M+m-1}{m}$ (number of multisets of size $m$ whose elements are taken from a set of size $M$). Now, let $f_{m}:\mathbb{R}^{m}\rightarrow\mathbb{R}$ be the $m$-th function in the given sequence of symmetric piecewise polynomial functions where each polynomial piece has degree at most $q$. Let $r$ be the number of pieces of $f_{m}$. Then, there is a subset of $\mathsf{Dom}(\Phi)$ with at least $\frac{1}{r}\binom{M+m-1}{m}$ elements where $f_{m}$ is a symmetric polynomial $P$ of degree at most $q$. Lemma LABEL:lemma:symmultipol tells us that $P$ can be expressed as a polynomial of degree at most $q$ of the elementary symmetric polynomials $s_{1},\cdots,s_{q}$. Due to the pigeonhole principle, any such polynomial will be equal on two distinct multisets $\{\{k_{1},\cdots,k_{m}\}\}$ and $\{\{k^{\prime}_{1},\cdots,k^{\prime}_{m}\}\}$ in $\Omega_{M}$ as soon as:
@@
Let $q$ be a positive integer. Let $F:\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}$ be a piecewise multivariate polynomial whose pieces are all $q$ generated. Let $G:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}^{d_{2}}$ be a function whose coordinates are piecewise polynomial functions whose pieces are all $q$-generated.
Then the pieces of the piecewise polynomial $F\circ G:\mathbb{R}^{d_{1}}\rightarrow\mathbb{R}$ are all $\tilde{q}$-generated, where $\tilde{q}$ depends only on $q$ and the degree of the pieces of $F$ and $G$.
@@
We first make the following observation. Let $A_{1},\cdots,A_{r_{1}}\subseteq\mathbb{R}^{d_{2}}$ be polynomial regions of $F$ and $B_{1},\cdots,B_{r_{2}}\subseteq\mathbb{R}^{d_{1}}$ be polynomial regions of $G$ that union to $\mathbb{R}^{d_{2}}$ and $\mathbb{R}^{d_{1}}$ respectively (such always exist, cf. Definition 1). For each $i\in[r_{1}]$, and $j\in[r_{2}]$, let $Z_{ij}:=(B_{i}\cap G^{-1}(A_{j}))$. Then:
  • $G$ and $F\circ G$ are polynomial on each $Z_{ij}$, as $Z_{ij}\subseteq B_{i}$, and $G(Z_{ij})\subseteq{A}_{j}$.
  • The $Z_{ij}$ union to $\mathbb{R}^{d_{1}}$, as $\bigcup_{i\in[r_{2}],j\in[r_{1}]}Z_{ij}=\bigcup_{j\in[r_{1}]}\bigcup_{i\in[r_{2}]}B_{i}\cap G^{-1}(A_{j})=\bigcup_{j\in[r_{1}]}\left(G^{-1}(A_{j})\bigcap\underbrace{\left(\bigcup_{i\in[r_{2}]}B_{i}\right)}_{\mathbb{R}^{d_{1}}}\right)=\bigcup_{j\in[r_{1}]}G^{-1}(A_{j})=\mathbb{R}^{d_{1}}$. The last equality follows from $\bigcup_{i\in[r_{1}]}A_{i}=\mathbb{R}^{d_{2}}\implies\bigcup_{i\in[r_{1}]}G^{-1}(A_{i})=\mathbb{R}^{d_{1}}$.
Hence, a polynomial piece of $F\circ G$ is of the form $F_{i}\circ G_{j}$ where $F_{1},\cdots,F_{r_{1}}$ are the polynomial pieces of $F$, and $G_{1},\cdots,G_{r_{2}}$ are the polynomials pieces of $G$.
If $F_{i}=\sum_{\alpha\in S_{F_{i}}}\gamma_{\alpha}X^{\alpha}$ and $G_{j}=\sum_{\beta\in S_{G_{j}}}\nu_{\beta}X^{\beta}$ with $\gamma_{\alpha}$ and $\beta_{\nu}$ integers then
$(F_{i}\circ G_{j})(X)$ $=\sum_{\alpha\in S_{F_{i}}}\gamma_{\alpha}\left(\sum_{\beta\in S_{G_{j}}}\nu_{\beta}X^{\beta}\right)^{\alpha}$
$=\sum_{\alpha\in S_{F_{i}}}\left(\sum_{k=1}^{q}\lambda_{k,\alpha}\sigma_{k}\right)\left(\sum_{\beta\in S_{G_{j}}}\left(\sum_{k=1}^{q}\eta_{k,\alpha}\sigma_{k}\right)X^{\beta}\right)^{\alpha}$
After expansion, the polynomial $(F_{i}\,\circ\,G_{j})$ has coefficients that are linear combinations of multivariate monomials of $\sigma_{1},\cdots,\sigma_{q}$ (of total degree at most the degree of $F_{j}+1$ in this case) with integer coefficients, as the $\lambda$ and $\eta$ coefficients are supposed to be integers. This shows that $F_{i}\circ G_{j}$ is $\tilde{q}$-generated, where $\tilde{q}$ depends only on $p$, the degree of $F_{i}$ and $G_{j}$. ∎
Lemma 3.
For positive integers $m$, $k_{1},\cdots,$ and $k_{m}$, let $T[k_{1},\cdots,k_{m}]$ designate the rooted tree illustrated in Figure 1. For any vertex $v\in V(T[k_{1},\cdots,k_{m}])$, Let $\xi^{t}(T[k_{1},\cdots,k_{m}],v)$ be the embedding of the vertex $v\in V$ obtained via a GNN with piecewise activation functions after $t$ iterations (where $\xi^{0}(u)=1$ for any vertex $u$ of $T[k_{1},\cdots,k_{m}]$). Then, for any iteration $t$, there exists an integer $q$ such that for any integer $m$, and any vertex $v\in V$, there exists a symmetric multivariate piecewise polynomial function $F_{m}$ such that
  • $\xi^{t}(T[k_{1},\ldots,k_{m}],v)=F_{m}(k_{1},\cdots,k_{m})$.
  • $\mathsf{deg}(F_{m})\leq q$
  • each piece of $F_{m}$ is $q$-generated
Proof.
We first prove all properties of $\xi^{t}(T[k_{1},\ldots,k_{m}],v)$ by induction on $t$.
Base case: for $t=0$ this is trivial since all vertices are initialised with the constant polynomial 1, whose degree does not depend on $m$, and is finitely generated.
Induction step: Suppose the property is true at iteration $t$, i.e there exists an integer $q_{t}$ such that for each vertex $w$, and for every integer $m$, $\xi^{t}(T[k_{1},\ldots,k_{m}],w)$ is a multivariate polynomial of the $k_{i}$’s whose degree is upperbounded by $q_{t}$, and is $q_{t}$ generated. For every vertex $v$ of $T[k_{1},\cdots,k_{m}]$